the incentre of the triangle formed by (1,2) (3,4) (2,3) is 


a)0  b)√2  c)1/√2  d)none


please explain

Grade:10

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Halak,

 

The incentre will be a co-ordinate. None of the option is in the co-ordinate format.

 

If it is inradus "r" that you are looking for,

Then r = area/semi-perimeter

 

Find a,b,c from distance formula. Then find area by Herons formula.

And s = (a+b+c)/2

And hence find the inradius.

 

You have a = 2root(2), b = root(2), and c = root(2).

 

Oh !!!, please note, the given points cannot represent a triangle, as they are collinear and lie on the same straight line.

 

So inradius = 0.

 

Hope this helps.

 

Best Regards,

Ashwin (IIT Madras).